Convert 43 lb to grams


1
Expert's answer
2019-09-19T09:39:28-0400

"43\text{ lb} = 43\text{ lb } \cdot \frac{0.45359237\text{ kg}}{\text{lb}} \cdot \frac{1000\text{ g}}{\text{kg}} = 19504.47191\text{ g} \approx 1.95\cdot 10^4 \text{ g}."

Reference: https://en.wikipedia.org/wiki/Pound_(mass)